Solve by using Elimination.
1/14x + 1/2y = 11/7
1/3x + 2/15y = -38/15

Answer by stanbon(75887)

Solve by using Elimination.
1/14x + 1/2y = 11/7
1/3x + 2/15y = -38/15
--------------------------
multiply thru the 1st equation by 14
multiply thru the 2nd equation by 15
-----------------------------------------
x + 7y = 22
5x + 2y = -38
--------------------
Multiply thru the 1st equation by 5:
5x + 35y = 110
5x + 2y = -38
---------------------
Subtract the 2nd from the 1st and solve for "y":
33y = 148
y = 148/33
---
Substitute into x + 7y = 22 to solve for "x":
x + 7(148/33) = 22
x = -310/33
--------------------------
Cheers,
Stan H.
================